OpenStudy (anonymous):

Use the percent proportion to find what number is 25% of 18

OpenStudy (jamesj):

25% of 18 is (25/100)x18 = 1/4 x 18 = 4 1/2 or 9/2

OpenStudy (anonymous):

Thank you !